How does the ideal mechanical


advantage (IMA) of an inclined plane vary


with the angle 𝜃 of the plane; that is, what


trigonometric function of 𝜃 gives the value


of IMA?

1
Expert's answer
2022-01-14T09:46:09-0500

"IMA=\dfrac{L}{H}=\dfrac{L}{L\sin\theta}=\dfrac 1{\sin\theta}=\csc\theta."

This trigonometric function is cosecans.